OpenStudy (anonymous):

sam has a total of 58 dvd's and cd's. if the number of cd's is two more than three times the number of dvd's how many cd's does he have A. 42 B. 14 C. 44 D. 12

OpenStudy (anonymous):

@PizzaLover123

OpenStudy (anonymous):

@west_coast

OpenStudy (clalgee):

Alrighty! Let's let d represent the amount of DVD's Sam has. Now, let's let c represent the amount of CD's he also has. Equation: c = 3d + 2 In the first line it tells us that Sam has a total of 58. Or if we add up the amount of CD's and DVD's Sam has then we will get 58. Look below: c + d = 58 Now add in the "two more than three times the number" 3d + 2 + d= 58 4d = 56 d = 14 c = 3 (14) + 2 = 42 + 2 = 44 So he has 14 DVDs and 44 CDs

OpenStudy (clalgee):

@anthony1798

OpenStudy (clalgee):

Also, welcome to Openstudy @anthony1798 !! \(\bf\Huge Welcome~to~\color{#00B4ff}{Open}\color{#7cc517}{Study!}\) \(\bf\Large\color{#00B4ff}{@anthony1798!!}\) \(\large\color{#7cc517}{\bf{-CLALGEE}}\)